OpenStudy (anonymous):

-100 + 97 -50 -23 + 57

OpenStudy (igreen):

Seriously..type it in google..

OpenStudy (anonymous):

-19

OpenStudy (igreen):

Just copy and paste it in Google..instant answer.

OpenStudy (anonymous):

-19

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

-19 IS THE CORRECT ANSWER

OpenStudy (muzzack):

-100+97=-3-50=-53-23=-76+57=-19

OpenStudy (anonymous):

you got 7 people on here lol

OpenStudy (dls):

Let, a=-100 b= 97 c= -50 d= -23 e= 57 the equation is a+b+c+d+e= Z (say) now you gotta use complex no.s and integrate to solve the equation :)

OpenStudy (anonymous):

YEAH LOL